OpenStudy (anonymous):

The height of the tunnel at the center is 27 ft and the vertical clearance must be 9 ft at a point 24 ft from the center. Find an equation for the ellipse.

OpenStudy (dumbcow):

 a = 24 b= 18 \[\frac{x^2}{24^2} + \frac{y^2}{18^2} = 1\]